OpenStudy (anonymous):

I will give a medal and become a fan if someone helps me. Henry opened a pack of paperclips. After using 8 of the paperclips, there were at least forty paperclips remaining. How many paperclips might have been in the box before Henry opened it? Let p represent the number of paperclips that might have been in the box before Henry opened it. p is greater than thirty-two p is greater than forty-eight p is greater than or equal to thirty-two p is greater than or equal to forty-eight

OpenStudy (joshawad):

40+8 is ___ approximately....

OpenStudy (anonymous):

p=paperclip total used=8 so left =p-8=40 p=40+8=48

OpenStudy (anonymous):

so p is greater than equal to 48

OpenStudy (anonymous):

thanks!!!! @g17